Suppose the real rate is 3.5 percent and the inflation rate is 5.1 percent.

  

What rate would you expect to see on a Treasury bill? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. (e.g., 32.16).)
Answer #1

r=.035
h=.051
R = (1 + r) * (1 + h) - 1
=(1 + .035) * (1 + .051) -1=0.0878=8.78%
Answer #2

ANSWER :



Rate on treasury bill 

= (1 + R)(1 +I) - 1 

= (1 + 0.035)(1 + 0.051) - 1

= 0.0878

= 8.78%  (ANSWER).

answered by: Tulsiram Garg